Skip to main content

IT Status Alerts

All Closed Alerts

Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250 251 252 253 254 255 256 257 258 259 260 261 262

5240 results found, displaying results 341 - 361.

Scheduled Service Change: Reboot OBIEE Prod server

Start:
2018-02-04 06:30:00
Status:
Closed
Brief Description:
Reboot the OBIEE production server to implement fixes.
more details

Scheduled Service Change: Blackboard maintenance - vendor fix

Start:
2018-02-04 05:00:00
Status:
Closed
Brief Description:
Blackboard vendor requires full system outage to fix critical error
more details

Scheduled Service Change: Move existing CMail users into OUs by their primary affiliation.

Start:
2018-02-04 08:00:00
Status:
Closed
Brief Description:
Move existing CMail/G-Suite users into OUs to make it possible to apply different policies to different populations.
more details

Scheduled Service Change: DNSDB prod server reboots

Start:
2018-02-01 04:00:00
Status:
Closed
Brief Description:
DNSDB prod servers( jonquil4 and sf-infranode052) will be rebooted on 2018/02/01 (Thursday) at 4am.
more details

Scheduled Service Change: PeopleSoft: Bi-weekly migration of online and batch objects 2/1/2018

Start:
2018-02-01 05:00:00
Status:
Closed
Brief Description:
PeopleSoft: Bi-weekly migration of online and batch objects. 2/1/2018
more details

Performance Issue: (Intermittent) Office 365 Email to Spectrum/Time Warner/RoadRunner

Start:
2017-12-01 12:22:00
Status:
Closed
Brief Description:
Cornell engineers have asked Microsoft to investigate intermittent refusal of email from Cornell’s Office 365 to an email account with Spectrum/Time Warner. Testing suggests that if a refused email is retried, it may eventually go through.
Current Status:
A ticket is open with Microsoft support.
more details

Scheduled Service Change: Net-Print PaperCut Prod Linux OS additional restart cycle

Start:
2018-01-30 04:00:00
Status:
Closed
Brief Description:
Net-Print PaperCut Prod Linux OS reboot cycle
more details

Scheduled Service Change: OS patching alternate e-list server (Y2)

Start:
2018-01-30 03:45:00
Status:
Closed
Brief Description:
The alternate e-list server (Y2) will be down for OS security patching
more details

Scheduled Service Change: ImageNow database server restart

Start:
2018-01-29 20:00:00
Status:
Closed
Brief Description:
Change Description: We will be bringing the system down so that the database server can be rebooted after OS patches are applied.
more details

Unplanned Outage: Password Manager Pro (PMP) Down

Start:
2018-01-29 08:45:00
Status:
Closed
Brief Description:
Password Manager Pro is experiencing an outage. Engineers in the IT Security Office are working to restore the system as quickly as possible. At this time, LastPass is unaffected.
Current Status:
Password Manager Pro is back up and fully functional. Please report any further outages to security-services@cornell.edu.
more details

Performance Issue: Thesis/Dissertation Approval System - Email Fraud Detection

Start:
2018-01-04 15:23:00
Status:
Closed
Brief Description:
The Thesis/Dissertation Approval System is currently experiencing an email issue related to it failing fraud detection checks.
Current Status:
We are currently working with both Media3 and the Email team to resolve the issue.
more details

Scheduled Service Change: EZ-Backup server is being rebooted to apply linux patching

Start:
2018-01-28 15:00:00
Status:
Closed
Brief Description:
EZ-Backup server is being rebooted to apply linux patching for Meltdown/Spectre
more details

Scheduled Service Change: SVC (San Volume Controller) maintenance in CCC and Rhodes Hall

Start:
2018-01-28 06:30:00
Status:
Closed
Brief Description:
Rescheduled from last week.
more details

Unplanned Outage: Wired and Wi-Fi data network outage at Rand Hall

Start:
2018-01-22 10:42:00
Status:
Closed
Brief Description:
The Rand1-5400 networking switch at Rand Hall, the foundry1-2530 networking switch and the hydroelec1-2530 networking switch are currently down.
Current Status:
These switches are now back up and wired and WiFi data connectivity has been restored to these buildings
more details

Scheduled Service Change: Net-Print PaperCut: Software Package Update

Start:
2018-01-21 07:00:00
Status:
Closed
Brief Description:
Net-Print PaperCut: Software Package Update bug fix.
more details

Scheduled Service Change: Remedy on Demand Maintenance - January 21, 2018

Start:
2018-01-21 00:00:00
Status:
Closed
Brief Description:
BMC (the vendor) will be performing regular monthly maintenance on the Remedy on Demand production systems Start: January 21 2018, 05:00 UTC (midnight EST) End: January 21s2018, 09:00 UTC (4 a.m. EST)
more details

Scheduled Service Change: Blackboard data center maintenance - impacting service use

Start:
2018-01-20 00:00:00
Status:
Closed
Brief Description:
On Saturday, January 20, 2018, Blackboard will be performing scheduled infrastructure maintenance.
more details

Scheduled Service Change: Relocate Fiber at Stimson Hall (Rescheduled to 1/20/2018)

Start:
2018-01-20 07:00:00
Status:
Closed
Brief Description:
Relocate fiber optic cable - RESCHEDULED from 1/13 due to weather to 1/20/2018.
more details

Performance Issue: Service restored, work ongoing:Verizon phones calling Cornell phone numbers

Start:
2018-01-18 08:54:00
Status:
Closed
Brief Description:
Verizon cell phones and Verizon landline phones were unable to call or receive calls from many (but not all) Cornell campus phone numbers. The caller may have heard a message indicating that the call did not go through.
Current Status:
AT&T has confirmed that this issue is resolved.
more details

Performance Issue: Issue with WhoIAm, netid.cornell.edu, applicantid.cornell.edu, etc.

Start:
2018-01-18 14:56:00
Status:
Closed
Brief Description:
WhoIAm, NetID Management (netid.cornell.edu), ApplicantID Management (applicantid.cornell.edu), Sponsored NetID creation, and new e-lists creation are affected by a server issue.
Current Status:
Service has been restored
more details
Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245 246 247 248 249 250 251 252 253 254 255 256 257 258 259 260 261 262